### v1

# Spica: Retrieving Scenarios for Pluralistic In-Context Alignment

Abstract
Alignment of large language models (LLMs) to societal values should account for pluralistic values from diverse groups. One technique uses in-context learning for inference-time alignment, but only considers similarity when drawing few-shot examples, not accounting for cross-group differences in value prioritization. We propose Spica, a framework for pluralistic alignment that accounts for group-level differences during in-context example retrieval. Spica introduces three designs to facilitate pluralistic alignment: scenario banks, group-informed metrics, and in-context alignment prompts. From an evaluation of Spica on an alignment task collecting inputs from four demographic groups ($n=544$), our metrics retrieve in-context examples that more closely match observed preferences, with the best prompt configuration using multiple contrastive responses to demonstrate examples. In an end-to-end evaluation ($n=80$), we observe that Spica-aligned models are higher rated than a baseline similarity-only retrieval approach, with groups seeing up to a +0.16 point improvement on a 5 point scale. Additionally, gains from Spica were more uniform, with all groups benefiting from alignment rather than only some. Finally, we find that while a group-agnostic approach can effectively align to aggregated values, it is not most suited for aligning to divergent groups. [^1] We provide our code and data for others to build on: https://github.com/Social-Futures-Lab/SPICA-code
1 Introduction
Figure 1: Example retrieval in traditional in-context alignment (ICA) systems rank examples based on similarity between prompts, failing to account for whether retrieved examples illustrate salient norms of a particular group. Spica addresses this limitation for pluralistic alignment by utilizing metrics to recover and incorporate each group’s own norms.
The widespread availability of generative AI systems has increasingly highlighted situations where outputs can be inappropriate, misleading, or dangerous to users Weidinger et al. (2021); Ji et al. (2023); Qi et al. (2024). Correspondingly, researchers have explored embedding human values into generative models through various value alignment strategies Huang et al. (2024); Gabriel (2020); Christian (2021); Ouyang et al. (2022). Since these models are typically deployed to the general public, model creators and service providers often seek to align towards a one-size-fits-all set of universal values Bai et al. (2022). However, different groups within society often disagree on values and have different norms around when and how to apply values Gordon et al. (2022); Weld et al. (2022); Park et al. (2024). Recognizing this challenge, more recent work has approached model alignment from a pluralistic perspective Sorensen et al. (2024b); Feng et al. (2024)—rather than attempt to bridge irreconcilable differences, a better solution may be to support different perspectives of each group directly.
One general strategy for large language model (LLM) alignment—in-context alignment (ICA) Lin et al. (2024); Han (2023)—acts dynamically at inference time by retrieving few-shot examples of prompts and associated preferable responses as context. ICA is a promising strategy for pluralistic value alignment as it flexibly allows the encoding of different groups’ values through swapping in and out each group’s own preferred responses. However, simply extending existing ICA to pluralistic preference data is insufficient—in practice, different groups differ not only on values, but also norms that dictate which values are more salient Weld et al. (2022). As current ICA retrieval metrics only consider relevance of the prompt and not the group, the in-context demonstrations can end up only illustrating less important values despite showcasing similar prompts (Figure 1).
In this work, we present Spica, an evolution of retrieval-based in-context alignment, enabling alignment to a plurality of group-specific values while being cognizant of group-specific norms around such values. Spica consists of three main components: (1) scenario banks—shared collections of scenarios (prompts, responses, and group preferences) that can encode both values and norms; (2) group-informed retrieval measures—metrics that allow us to recover second-order norms from individual preference assessments; (3) ICL prompt setups that can effectively apply richer information from scenarios to the task of alignment.
We evaluated Spica by conducting an alignment task for four demographicgroups and examined three main aspects of the process: the quality of the scenarios retrieved, the effectiveness of different in-context prompts in applying scenarios to alignment, and the performance on the end-to-end task of alignment of model outputs.
In our evaluation, we find that:
  • Compared to a baseline using only similarity-based scoring of examples, group-informed metrics retrieved scenarios that aligned more accurately (as measured by RMSE of ratings) to observed ground truth, indicating a quality gap caused by the reliance on similarity.
  • Among different ICL setups for integrating scenarios into prompts, the most effective designs were: P-I style—provide a single positive instruction when user preferences are collected over descriptions of response strategies; and C-R style—provide a contrasting spectrum of example responses when user preferences are collected over model outputs.
  • In an end-to-end evaluation, we find that Spica improves alignment overall more than baseline ICA (+0.053 / 5 points), with statistically significant gains (+0.16 / 5 points) observed on groups traditionally disadvantaged in alignment.
  • We also find that baseline ICA can result in disparate outcomes, with some groups preferring original non-aligned model outputs, whereas Spica alignment produces outputs uniformly preferred by all groups.
  • Finally, on alignment tasks for aggregated (collective) values, existing group-agnostic approaches can still perform well, despite not generalizing to pluralistic settings.
2 Related Work
Customizing LLMs for Value Alignment
Traditional methods for customizing LLMs for specific tasks and domains involve modifying training procedures. These include pretraining on task-specific corpora Wu et al. (2023); Lee et al. (2020), post-hoc finetuning Gururangan et al. (2020); Han and Eisenstein (2019), instruction tuning Ge et al. (2023); Gupta et al. (2022); Shi et al. (2023), and aligning with human preferences Ouyang et al. (2022). These approaches are also used to encode moral values and diverse human preferences into models Tay et al. (2020); Bai et al. (2022); Liu et al. (2022); Bang et al. (2023); Jang et al. (2023). However, they have significant limitations for value alignment. They require extensive human annotation to provide meaningful signals about desired values Kim et al. (2023), and even then, there is limited understanding or guarantee of how well the models have internalized these values Agarwal et al. (2024). This makes the models less robust in terms of value alignment. Moreover, once trained, these models lack flexibility; updating the model to reflect evolving values often requires a complete retraining, which is computationally intensive Carroll et al. (2024).
Retrieval-based In-Context Learning for Alignment
In-Context Learning (ICL) offers promising alternatives for value alignment by enabling behavior modifications during inference rather than training through the use of few-shot examples incorporated into model prompts Dong et al. (2022); Wei et al. (2022). The use of example demonstrations in ICL has also allowed systems to incorporate retrieval Lewis et al. (2020); Borgeaud et al. (2022) as a part of dynamically constructing in-context prompts informed by inputs Zhang et al. (2022); Rubin et al. (2021).
For the task of model alignment, approaches to using retrieval and in-context learning prompts, such as URIAL Lin et al. (2024), have also been referred to as in-context alignment (ICA) Han (2023). As most ICA systems today focus on addressing more collective preferences, the retrieval aspect of such systems has largely remained unchanged, with relatedness metrics like semantic similarity being the main way to rank retrieved examples Karpukhin et al. (2020); Gao et al. (2023). However, given examinations of how models apply in-context examples Min et al. (2022), there is room for improvement around how examples are ranked for retrieval, especially for value alignment tasks. Prior works around alignment have pointed to potentially important factors that could improve the utility of retrieved examples, such as noting potential importance for prioritizing examples that illustrate exceptional circumstances and edge cases Kiehne et al. (2022), or emphasizing examples that capturing population-specific preferences Hovy and Yang (2021); Kirk et al. (2023), with potentially more signals to consider for pluralistic alignment, where different groups can have different norms Weld et al. (2022) that moderate how preferences are prioritized over each other.
Accounting for Pluralism in Value Alignment
Supporting pluralistic values is crucial for building general-purpose agents and LLMs Sorensen et al. (2024b) that are deployed into society. Large datasets like ValuePrism Sorensen et al. (2024a) and PRISM Kirk et al. (2024) highlight the importance of reflecting diverse values, yet achieving consensus remains challenging. Some approaches turn to higher-level abstract descriptions of values as a solution for improving consensus Bai et al. (2022), however, practical application to specific cases often reveals discrepancies Koshy et al. (2023). Drawing from the legal realm, there have also been approaches that propose combining higher-level descriptions with specific examples (e.g., legal precedents) to illustrate more ambiguous concepts encoded by values Cheong et al. (2024); Chen and Zhang (2023).
Beyond first-order challenges of encoding values themselves as mentioned above, pluralism can also give rise to second-order challenges when groups share similar sets of preferences or values (such as preferring diversity and factual quality) while also disagreeing on their salience Jackson (1960) (and thus prioritization) when it comes to practical application Weld et al. (2022). This aspect, however, is often overlooked by existing frameworks for pluralistic alignment.
3 Retrieving Scenarios for Pluralistic In-Context Alignment (Spica)
Figure 2: Diagram illustrating the main components of Spica: (A) Collections of prompts, responses, and individual preferences form scenario banks which ground the alignment process; (B) During the ICA retrieval process, we make use of group-informed metrics to recover group values and norms, together with semantic similarity, these scores guide the ranking of scenarios; (C) Retrieved scenarios are incorporated into ICL prompts that make use of preference distributions and example responses to form alignment demonstrations.
In this work, we outline Spica, a framework that builds on existing ICA but with a specific focus on retrieving Scenarios for Pluralistic In-Context Alignment. Following this section, we will present three novel components of Spica (Figure 2), addressing: (1) how to encode group-specific values and norms in the form of scenario banks; (2) how to utilize the encoded group-specific norms during the retrieval process through group-informed metrics; and (3) how to make use of more nuanced preferences as encoded by scenarios through alternative designs for in-context learning prompts.
3.1 Scenario Banks for Encoding Pluralistic Values and Norms
Past examples of data for alignment has included both normative data in the form of guidelines created through constitutional alignment processes Bai et al. (2022) and quantitative data in the form of user ratings of conversations between humans and LLMs Kirk et al. (2024). While this data can be incorporated in ICA settings via prompts and example demonstrations respectively, it can be insufficient for the needs of pluralistic alignment. It can be difficult to attribute normative guidelines created through deliberation and consensus to each possible group in a population. Additionally, human ratings of conversations often only apply to existing outputs making it difficult to evaluate the broader space of responses.
Given this, with Spica, we propose a way of collecting pluralistic alignment data in the form of scenario banks. A scenario consists of three main components: (1) a prompt ($x$), which represents an example of a user query or conversation with a model leading up to a morel response; (2) responses ($y\in Y_{x}$), the space of possible ways a model could respond to a prompt, which can take the form of either specific examples of outputs, or high-level strategies covering a class of outputs; and (3) preferences ($r_{p}(x,y)$), which encode an individual $p$’s preference of a response $y$ to a prompt $x$. A scenario bank consists of a collection of such scenarios and provides a basis for the ground truth in an ICA retrieval. By creating scenario banks this way, we are able to understand distributional properties of preferences both at an individual level and at a group level, allowing us to understand both individual values and group norms in pluralistic alignment.
First, encoding values as distributions over responses provides better understanding to human preferences in context. Traditional human ratings collected for disjoint prompt-response pairs lead to rating values that need to be compared between prompts—e.g., there is not sufficient information to tell if a particular low-rated response to a prompt was just a bad response, or actually the “best possible” out of a prompt with no good options. Even when multiple responses from different models are evaluated against each other Kirk et al. (2024), we are still limited by distributional properties of the pre-alignment outputs which could lead to groups less represented by the training data struggling to indicate desirable outputs to them.
Second, by focusing on individuals, we can recover norms by looking at distributional properties within groups—even when groups may be (re-)constructed post-hoc. Theory around the structural nature of social norms Jackson (1960) pose that norms can be fundamentally observed as patterns in preferences that are salient across individuals within a group, which can now also be measured through scenario banks. Additionally, while other normative approaches like directly asking for high-level criteria can also uncover norms, outputs can be hard to disentangle as individuals can often have intersectional identities that span multiple groups. With scenario banks, one only needs to find the individuals corresponding to a group should group criteria change or new groups be considered, adding new data only if insufficient.
Finally, distributional data allows us better understanding of alignment outcomes, which can be especially important due to the stochastic nature of outputs. Given a ground truth rating distribution (assuming discrete response classes) $r(x,y),\forall y\in Y_{x}$, we can compare different alignments by simply observing the amount of divergence between the two. In this discrete case, we can simply use a loss-style function like the root mean squared error (RMSE):
$L(r(x),r^{\prime}(x))=(\sum_{y\in Y_{x}}(r(x,y)-r^{\prime}(x,y))^{2})^{\frac{1}{2}}$ (1)
We can use a similar approach to understand retrieval in ICA, where we can find the expected loss across a group if the model produced outputs similar to the retrieved $x^{\prime}$ but for $x$: $\mathbb{E}_{r}[L(r(x^{\prime}),r(x))]$.
3.2 Group-Informed Retrieval Measures
As noted earlier, in pluralistic alignment, values alone are often insufficient as communities Weld et al. (2022) and demographic groups Kumar et al. (2021) can have similar preferences on specific examples while making different higher-level trade-offs around what are salient examples—e.g., across many examples one may find that a group is prioritizing correctness over respectfulness, or helpfulness over safety, even when each property assessed individually comes out as desirable. Existing retrieval metrics which only compare similarity of the input prompts $x$ and known examples $x^{\prime}$ have no way to encode this group-level difference.
For a solution to this, we take inspiration from prior work that examines structural properties of social norms, with one such example being the return potential model Jackson (1960). This model interprets norms as distributional properties over behavior appropriateness that arise when viewing a group in aggregate. In the alignment setting, we adapt two concepts to assess model outputs (behavior): (1) We adapt the idea of “crystallization”—whether preferences over a behavior is consistent across group members—into the metric $g_{\text{stability}}(x^{\prime})$; and (2) we adapt the ideas of “intensity” and “tolerable range”—which assess differences in preference across behaviors for the same group member—into the metric $g_{\text{contrast}}(x^{\prime})$.
3.2.1 Stability: Differentiating Common Values from Individual Preferences
With social norms, “crystallization” assesses whether preferences on any particular behavior has become shared among the group, and thus become more crystallized as a norm. In the ICA setting, this concept can also be useful: While all ICL examples have the potential to illustrate preferences within a group, an example that illustrates a set of shared (crystallized) preferences is likely to be a better example demonstration for the group. We measure this for each scenario in our scenario bank: if, for that scenario $x^{\prime}$, the variance between annotators’ preferences $r_{p}(x^{\prime},y^{\prime})$ on each response type $y^{\prime}\in Y_{x^{\prime}}$ is lower, then the scenario is likely to demonstrate more crystallized norms than weaker preferences. More concretely:
$\text{stability}(x^{\prime},y^{\prime})$ $=-\frac{\sum_{r_{p}}(r_{p}(x^{\prime},y^{\prime})-\bar{\text{r}}(x^{\prime},y^{\prime}))^{2}}{|\{r_{p}\}|}$ (2)
$g_{\text{stability}}(x^{\prime})$ $=\mathbb{E}_{y^{\prime}}\left[\text{stability}(x^{\prime},y^{\prime})\right]$ (3)
3.2.2 Contrast: Assessing Indifference versus Preference
With social norms, concepts like “tolerable range” and“intensity” assess how broad the range of acceptable (and unacceptable) behaviors is and the intensity at which individuals express this preference Hackman (1992). Like with crystallization, we can also assess a similar concept with scenario banks: For a scenario $x^{\prime}$, the variance between different behaviors $y^{\prime}\in Y_{x^{\prime}}$ across each annotator $r_{p}(x^{\prime},y^{\prime})$ assesses how much they care about differentiating preferences. More concretely:
$\text{contrast}(x^{\prime},r_{p})$ $=\frac{\sum_{y^{\prime}}(r_{p}(x^{\prime},y^{\prime})-\bar{\text{r}}(x^{\prime},y^{\prime}))^{2}}{|\{(x^{\prime},y^{\prime})\}|}$ (4)
$g_{contrast}(x^{\prime})$ $=\mathbb{E}_{r_{p}}\left[\text{contrast}(x^{\prime},r_{p})\right]$ (5)
3.2.3 Learning Metric Weights
As our new metrics don’t replace relevance but only add group-level information, we need to balance potentially competing objectives. In Spica, we do this by taking a linear weighted combination of the introduced metrics and similarity (distance), in the form of: $\bar{d}(x,x^{\prime})=w_{d}\cdot d(x,x^{\prime})+w_{s}\cdot g_{\text{stability}}(x^{\prime})+w_{c}\cdot g_{\text{contrast}}(x^{\prime})+c$.
As the optimal weighting of metrics likely to vary across groups, we need to empirically find these weights. As discussed in Section 3.1, the retrieval desirability of $x^{\prime}$ given input $x$ can be assessed as an expected loss $\mathbb{E}[L(r(x),r(x^{\prime}))]$, we can compute the loss and minimize using a simple linear regression $\bar{d}(x,x^{\prime})=\mathbb{E}[L(r(x),r(x^{\prime}))]$. Here we show the case where only the best ($k=1$) is selected, though a similar approach could be applied for selecting a larger set of ${x^{\prime}}$.
3.3 In-Context Learning Prompts for Retrieved Scenarios
Because retrieved scenarios contain preference distributions across multiple responses (or strategies), different setups for integrating scenarios as demonstrations are likely to produce different model outputs. ICL prompt designs have been extensively studied by prior works, so in this work we primarily explore new configurations enabled by our construction of scenarios in the scenario bank.
Specifically, we focus on two factors: (1) the way preference distributions are incorporated into the prompt (The highest Positive preference or Contrasting pairs of highest and lowest preferences), and (2) the type of response feedback used (Instructions or example Response text). This results in 4 combinations of prompt setups: P-I, C-I, P-R, and C-R. Through human evaluations, we assess these prompting setups, with specific details of how we implement these prompts documented in Appendix A.6.2.
4 Experiments and Results
To evaluate Spica, we set up a small-scale proof-of-concept pluralistic alignment task involving 4 demographically constructed groups, and assess how well a Spica workflow is able to align model outputs to preferences of each group compared to a baseline approach that only considers semantic similarity. We conduct this evaluation over two language models and explore variations of the prompt used to incorporate scenarios. We assess the end-to-end alignment outputs and also examine the impact of pluralistic versus collective alignment.
4.1 Dataset and Scenario Bank Construction
For our evaluation alignment task, we constructed a set of queries (which define the topics to provide alignment on) by drawing from an existing set of challenging alignment situations based on prompts observed in conversations on the PRISM dataset Kirk et al. (2024). PRISM engaged human participants to interact with LLMs by naturally starting conversations with 3 types of guidance meant to invoke conversations around more challenging and complex topics: “unguided”, “values guided”, or “controversy guided”. We observed that of the 3 types of guidance, unguided conversations primarily resulted in simple informational requests which are not particularly controversial in the context of pluralistic alignment, so we opted to drop conversations of this type. Among the remaining conversations, we randomly selected a subset, split into 3 slices: retrieval (train, $n=360$), weight optimization and prompt strategy evaluation (dev, $n=150$), and evaluation hold-out (test, $n=75$).
For the responses used in our scenario bank, we did not use responses from PRISM, instead we follow Section 3.1 and construct responses for each aligned model ourselves based on a set of strategies covering common behaviors Appendix A.6.1). To capture the stochastic nature of model outputs, we generate 3 responses for each strategy.
4.2 Models and Similarity Metric
For our experiments, we tested the quality of retrieval-based ICL alignment using one open-source (llama3-8b) and one closed source model (gpt-4o-2024-05-13) as the base model. llama3-8b inference was conducted using a locally hosted instance of Ollama  [^2] https://ollama.com/. With both models, we applied the same prompts to generate responses attached to scenario bank queries and to conduct apply in-context alignment (Appendix A.6.2).
As our goal is to evaluate the additional metrics we introduced, we held the semantic similarity measurements constant across all models and conditions, using values derived by computing the cosine similarity between embeddings generated by text-embedding-3-large.
4.3 Pluralistic Groups and Human Annotation Setup
For our pluralistic alignment task, we define four groups in the form of demographic slices drawn from the US populations: partisan political affiliation (“Republican” or “Democrat”), and self-reported regular participation in religious activities (“yes” or “no”). Our choice of these specific features is based on these factors being salient for different opinions around AI Zhang and Dafoe (2019) along with practical considerations around available demographics recruitable on our crowd platform Prolific.
Annotators in each group participated in providing preference assessments over our dataset, in the form of an annotation survey (Appendix 11) where they were shown 15 prompts from the dataset, each of which included 1 response for each of the 5 strategies. Participants rated both the output and the strategy associated with the output in terms of appropriateness (on a scale 1 - 5). Combined with 5 attention checks, participants completed a total of 80 sub-tasks with a median time of 30 minutes. For the annotation portion, we recruited a total of 544 participants to cover the annotation on train and dev sets across two model types, guaranteeing 2 annotations per group per scenario.
For the end-to-end evaluation, we recruited additional annotators from each group, who assessed outputs produced after ICL alignment using a similar rating survey interface to the annotation process, instead framed as rating outputs from different systems rather than different strategies. For each end-to-end evaluation, in addition to participants who evaluated outputs aligned to their own group, we also set aside 1/3 of the users from each group to evaluate the outputs of alignment against all 4 groups aggregated as one large group to simulate a more collective value alignment setting (Section 4.7). We recruited a total of 240 participants to conduct the evaluation of prompting strategies (Section 4.5) and 120 participants for the end-to-end evaluation on the held-out test set (Section 4.6). Tasks were calibrated on Prolific to pay at a rate of \$12 USD/hour with the study design was reviewed and determined exempt by our IRB.
4.4 Results: Evaluating Retrieved Scenarios
Figure 3: Average error over preferences (measured by RMSE) comparing retrieved scenarios and observed ground preferences on the dev set. baseline uses similarity-only retrieval. spica uses the weighted group-informed metrics.
For our first evaluation, we examine whether group-informed metrics can improve the scenarios retrieved. To evaluate this, we look to the average error over preferences (RMSE) metric introduced in 3.1: A scenario with a lower error means that, for each response type in the scenario, preferences about its appropriateness are more similar to the actual observed preferences on that type of response for the input query. This evaluation is possible as we use queries from scenarios in the dev set, where preferences have also been collected over each response type. We follow the design of Spica to retrieve scenarios from the scenario bank (train) as ICL input examples. For the baseline condition, we only use the similarity metric over the prompts, while with Spica we utilize the additional group-informed metrics over preferences.
As indicated in Figure 3, we see that for both models, Spica is able to retrieve scenarios that had preferences more accurately matching those observed directly on the dev set. We also examined this effect at a per-group level, and find that while error is lowered across all groups, the magnitude of such change can differ (details of this we examine in Appendix A.1).
4.5 Results: Evaluating In-Context Prompting Strategies
Figure 4: Comparison of end-to-end human evaluations of alignment outputs on the dev produced by the 4 prompting setup combinations: Positive-only or Contrastive, Instructions or example Responses.
In order to examine the effectiveness of ICL prompting setups (Section 3.3), we used human participants to evaluate the outputs produced models given each type of prompt while using the same Spica retrieval setup. Participants evaluated the outputs using an interface similar to that used during preference collection for scenario banks, however, instead of rating response strategies, participants rated 5 hypothetical AI systems (“System A - E”), each representing one configuration with a final control output produced by the model with no ICL alignment. Results are evaluated within-subjects, computing the difference between the rating of each condition’s aligned output and the control output, presented as “rating delta”.
As shown in Figure 4, we find that, for the gpt-4o model in a pluralistic alignment setting, the combination of contrastive response examples (C-R) proved to be the most effective (significant $p=0.030<0.05$ via ANOVA), on average rating $0.10$ points higher than the control across all groups. Additionally, given that scenario bank preferences may be collected only over high-level strategies, we also find positive instructions (P-I) to be the most effective prompt setup in this situation, producing ratings that average $0.07$ points higher (though not statistically significant). Given the results from this evaluation, the most effective prompting setup as observed on gpt-4o are P-I) and C-R. We use these two configurations in our final end-to-end evaluation on the test set, referring to the Spica with these prompt setups as Spica-I and Spica-R respectively.
We also examined these ICL prompts on the llama3-8b model. However, here none of the setups provided reliable alignment outcomes. In the pluralistic alignment task, we found no statistically significant differences between outputs, and average change in ratings were small or negative. We hypothesize that the smaller size of the llama3-8b model we used may be less capable at generalizing from ICL-style alignment.
4.6 Results: Evaluating End-to-End Alignment Outputs
Figure 5: End-to-end human evaluation of group-aligned outputs on the test set user queries for both models. Figure presents the aggregated results across the 4 group alignments.
Given the above findings on configurations, we conducted end-to-end evaluation that generate outputs for a held out test set of user queries. For comparison, we test a baseline ICA setup, where retrieval only considers semantic similarity with ICL prompt incorporating the highest rated response in each scenario as the demonstration example. For Spica, we use the two prompting strategies from above: Spica-I and Spica-R.
In Figure 5, we present results for the final end-to-end evaluation. We observe that on llama3-8b, none of the ICL alignment approaches performed well compared to the unaligned control, with the highest Spica-R averaging +0.005 / 5 rating points above control and no statistically significant differences between conditions overall or at a group-level. On gpt-4o, ICL alignment is effective, with the best system Spica-R performing better +0.072 / 5 points better than the unaligned control.
While considering all groups collectively, no condition was significantly better, if we look at each group individually, we find that for the rep-nrel group, Spica-R did result in statistically significant +0.16 points higher performance compared to baseline (within subjects paired t-test, $p=0.044<0.05$), with the rep-rel group also seeing a borderline significant improvement (within subjects paired t-test, $p=0.051$) of +0.16 points. Considering that recent work Rozado (2024) has found many current LLMs to favor liberal values, this result suggests that pluralistic alignment via Spica provides benefits primarily by improving alignment for traditionally disadvantaged groups.
Further examining alignment outcomes at a group level (Figure 9) we also find support that Spica can lead to more equitable outcomes across groups: With baseline on gpt-4o, we find that while the dem-rel and dem-nrel groups prefer aligned outputs (seen as +0.11, and +0.13 points over control), the rep-rel and rep-nrel groups end up preferring the original outputs (observed as -0.07, and -0.11 rating points compared to control). This discrepancy between groups is statistically significant for the minority group of non-religious Republican identifying (rep-nrel) participants (unpaired t-test between groups, $p=0.031$ and $p=0.049$). However, with Spica-R, all groups now prefer aligned aligned outputs (+0.10, +0.05, +0.09, +0.05) and we no longer see any statistically significant difference between groups in terms of this preference. Considering that in all conditions, ICL examples draw from each group’s own preference annotations (values), this supports the importance of considering group characteristics (norms) as a part of the retrieval process.
4.7 Results: Comparing Pluralistic versus Collective Alignment
One final aspect we wanted to explore at the end, was whether similar characteristics from pluralistic alignment carry over to a traditional collective value alignment setting. To accomplish this, in the evaluations for Section 4.4 and Section 4.6, we also included an additional type of system output—outputs aligned against all—where we used the same setup for each condition, with the only difference being that instead of having group-specific preferences, we define an “all” group that aggregates annotated preference judgments from all 4 groups. During evaluation, a portion of participants (Section 4.3) from each group would then be given this collectively aligned output to rate instead of group-specific alignment outputs.
In Figure 6, we see that unlike the pluralistic setting, when aligning to collective values of this super-group, optimal ICL prompt strategies on gpt-4o now significantly favor instructions (P-I and C-I) over example responses (llama3-8b still cannot be effectively aligned with either).
On the end-to-end evaluation of test set queries, while Spica-R is no longer an optimal prompt style for gpt-4o, we do see that group-informed retrieval metrics from Spica also no longer seem to provide any benefit, even slightly under-performing baseline retrieval. This is unsurprising though, as we would not expect pluralistic group-informed metrics to contribute significantly to the selection of scenarios for a “group” with few shared norms. Instead, now we find that the ICL prompt style becomes the most important factor, with gpt-4o favoring instructions and llama3-8b favoring example responses (used by both baseline and Spica-R), with this trend largely carrying over even across the participant’s own group affinity differences. This suggests that for traditional collective alignment, baseline retrieval setups are sufficient.
(a) Evaluating different ICL prompt setups on the all group over the dev set scenarios.
(b) Evaluating end-to-end outputs all group over the test set queries.
Figure 6: Results of the same evaluations as used in Section 4.5 and Section 4.6, but only defining one group (all) that aggregates the preferences of all 4 of our pluralistic groups.
5 Discussion
5.1 The Role of Pluralistic Versus Collective Value Alignment
Considering the results of our experiments, there is a discussion to be had around the role of pluralistic versus collective value alignment. From prior work, we have seen how existing models can favor the values and norms of their designers and of majority populations Buyl et al. (2024); Rozado (2024), and how traditional alignment setups can be insufficient when addressing the challenges introduced by pluralistic alignment. However, while the groups in our evaluation were selected to illustrate how norms and values can differ, it is not hard to imagine how pushing for pluralistic alignment along demographic and ideological lines could lead to seemingly happier groups and communities at the cost of increased social fragmentation. Indeed, processes for consensus during collective value alignment Bai et al. (2022) can be means for society to find or build common values across groups.
Ultimately, we see a role for both pluralistic and collective value alignment in the broader landscape for building better socially-situated LLMs. Building from measurements around social norms like those used in Spica, there are opportunities for identifying pluralistic groups that reflect healthy diversity within society rather than simply being divisive. At the same time, collective value alignment processes, especially ones involving democratic inputs, remain important tools to bridge divides and find common ground.
5.2 Towards More Efficient Ways to Map of Group Values and Norms
In this work, we constructed scenario bank scenarios by drawing prompt data from an existing alignment dataset. While this can be a simple way to bootstrap the Spica framework, it is also not particularly efficient when it comes to mapping out meaningful norms and values of a group. In practice, we found that scenarios created in this way often end up illustrating similar values, leading to duplicated efforts during the annotation process for limited benefit. One avenue for improving the efficiency around eliciting group values that has been proposed by prior work is to engage group members in actively probing their own values through human-in-the-loop interactive systems Klingefjord et al. (2024). While current systems have primarily focused on more abstract value statements, similar designs could be adopted for groups to interactively expand upon scenario banks guided by their own exploration of the space of scenarios relevant to them. We believe there are many opportunities here for both new interactive system designs around collaboratively synthesizing scenarios, as well as new metrics that dynamically surface regions of scenarios that have been overlooked by each group.
5.3 Generalizing to Non-Discrete Preferences
In the specific implementation presented in this work, we apply Spica primarily in a discrete setting, defining preference functions through rating scores and responses through either concrete text or high-level strategies. While we discuss these limitations, we do also note that in practice alignment often applies to continuous settings: prompts, responses, and preferences all lie in a continuum.
While we do not explore this, the ideas introduced in Spica can largely generalize to non-discrete settings with few modifications. For example, while prompts and responses are currently sampled discretely, a non-discrete version of Spica may make use of embeddings and diffusion-based generative models Li et al. (2022) to examine regions of a continuous generation space. The measures in Spica could also be applied in settings where continuous preferences can be acquired, with concepts like “stability” and “contrast” potentially being captured through distribution-level properties like divergence and kurtosis.
Limitations
External Safeguards
While this work explores in-context learning approaches to value alignment, the models we use as a source to build aligned models from also come with their own existing safeguards. In the case of gpt-4o, this can be especially challenging, as the model itself is not open source and only being able to interact with such models through a proprietary API likely leaves the possibility of external safeguards. These safeguards are often implemented as additional models that may monitor the inputs and outputs around the main LLM, jumping in to intervene should it identify inappropriate inputs being supplied or outputs being generated. This means our ability to affect the outputs of such models may be limited in ways that cannot be addressed by modifying the prompts.
Strategy Adherence
In our study, we use a set of 5 response strategies to approximate a diverse set of responses for each prompt. While there is evidence from prior work that human preferences tend to align towards high-level strategies Cheong et al. (2024), generating responses following fixed strategies may not always be reliable as responses may not always adhere to the strategies (either due to model safeguards or relevance of the strategy to an input prompt). To control for the effects of this, during our annotations of the scenario bank, we asked annotators for input on both concrete responses and high-level instructions and only used the corresponding rating data when testing prompting strategies based on instructions versus examples. Still, this may be insufficient to address the resulting reduction in variation of the response space on some prompts.
Participants and Scale
In our experiments, we focused primarily on a small-scale proof-of-concept alignment task targeted towards a US population. As a result, we were only able to examine the outcomes of alignment over one source of input prompts (PRISM) and several demographically-constructed groups based on US participants. While in this setup, we did observe differences between alignment mechanisms and also differences between goals (pluralistic alignment at a group level versus alignment towards the entire population), it is likely that given different constructions of the groups, there may be different takeaways.
Ethics Statement
The AI alignment problem itself has many ethical implications, and these considerations also extend to both implications of the the design of Spica, and our choices during our evaluation of it.
First, our experiments are intended to demonstrate a proof-of-concept setting where different groups are likely to have significant divergent values, as a result of this consideration and practicalities surrounding ease of recruitment, we we opted to extrinsically define “groups” based on divisive demographic features within a US-based participant pool. However, this should not be interpreted as an endorsement for using politics and religion as a way to conduct pluralistic alignment—many other factors like culture, community, and identity could provide better delineation between different groups with lower risks around introducing additional social fragmentation. Given this, we also caution against using results in this work to make inferences about the broader population groups we tested with, as we didn’t make additional efforts to ensure our participants are representative samples within these groups.
Secondly, to emphasize how values can differ, we drew our evaluation scenarios from the PRISM alignment dataset in a way that prioritizes controversial scenarios (Section 4.1). Coupled with limitations in PRISM’s data collection itself, it is likely that the distribution of scenarios would be biased towards being able to better capture certain values over others. The goal of our setup is to ensure potential biases of this sort at least are applying to all tested conditions, so we also caution against using our results to make inferences about the alignment scenarios themselves.
Finally, there are ethical considerations around the basic motivation for pluralistic alignment Jiang et al. (2024). By allowing groups and communities to build AI tools that reflect their own values, we run the risk of producing self-reinforcing echo chambers; thus, while we don’t focus on aspects beyond social preferences, we do recognize that other aspects of alignment (factuality, diversity, fluency, etc.) remain important problems that cannot be addressed by frameworks like Spica as-is.
Appendix A Appendix
A.1 Results: Group-level Breakdown of the Retrieval Loss
Figure 7: Group-by-group breakdown of the difference in retrieval quality between baseline semantic similarity and Spica.
We present a group-by-group breakdown of the retrieval loss in Figure 7. Interestingly, we find that the groups indicting higher affinity to religion (-rel) tended to see a more marked difference in retrieval quality. This seems to be the result of these groups having more preferences over responses that are not as dependent on the specific prompt and instead apply to a wide variety of topics. For gpt-4o, the P-I and C-R conditions consistently produced positive alignment outcomes.
A.2 Results: Group-level Breakdown of Prompt Strategy Results
Figure 8: Group-by-group breakdown showing differences between groups in their evaluation of outputs produced though different prompts on the same retrieved examples.
We present a group-by-group breakdown of the prompting strategy evaluation in Figure 8. Interestingly, we note that while there are some consistent trends (such as only using a single positive example for example responses), prompt strategy effectiveness can also vary significantly across different population groups. For example, contrasting prompts worked well for aligning preferences for the rep-rel group, while instruction-based prompts worked well for the rep-nrel group. While this should not be seen as generalizable takeaways for properties of specific populations, it is still important to note that ICL prompting strategy effectiveness can vary depending on the group (or, more relevantly, the norms and values exhibited by the group).
A.3 Results: Group-level Breakdown of End-to-End Evaluation
Figure 9: Group-by-group breakdown showing differences between groups in their evaluation of outputs on the final end-to-end task. Green indicates Spica-retrieval + prompting based on presenting instructions for the best response strategy of the retrieved instances. Blue indicates Spica-retrieval + prompting based on showing contrastive example responses associated with the retrieved instances.
We present a group-by-group breakdown of the final end-to-end evaluation in Figure 9. For gpt-4o, we found Spica with contrastive examples to be the provide the most consistent alignment across groups, being preferred over the control response, but not always the most preferred response across the alignment conditions. Baseline retrieval seemed to be effective in alignment for dem-identifying groups but produced the opposite outcome for rep-identifying ones.
A.4 Results: Qualitative Analysis of the Weights Learned
Figure 10: Final weights learned for each group and alignment target learned from the dev set. The Spica composite metric represents a distance (in this case modeled by the loss), which we want to minimize. Metrics represent scores, with higher values indicating more, hence the coefficients are primarily negative. strategy-rating indicates values produced by using user ratings over response types, while response-rating indicates values produced by user ratings over response examples.
Finally, we qualitatively look at the weights learned for various groups across each model. Here we observe that, weights produced after learning from response types preferences and response example preferences end up relatively similar to each other. We also note that similarity scores (in this case cosine similarity) receives a comparatively lower absolute weight compared to the other metrics, however, this is as expected, as similarity scores tend to span a different range of values than preference level metrics. We also observe that between the two new metrics, stability is the most important for the all experiment, matching the notion that in a collective alignment setting, using examples that are closer to universal values tends to be more ideal, while at a group-level there is no such pattern. Finally, for the -nrel groups we observed cases where similarity was assigned a positive weight, implying that examples immediately closer to the query were actually often less desirable, possibly a reflection of non-religious groups finding subject matter around different religious topics less similar to each other than religious identifying groups. Beyond this, though, the weights seem overall unsurprising, with no other significant patterns of note.
A.5 Human Annotation Materials
In this section we document the instructions and materials used for our human annotation and evaluation tasks.
A.5.1 Instructions
We are researchers from [REDACTED] and we are conducting a study to understand people’s preferences on the behavior of generative AI chatbots or virtual assistants. Generative AI chatbots and assistants (examples include OpenAI’s ChatGPT, Microsoft Copilot, and Google Gemini) are computer programs designed to generate text in response to user questions or prompts. However, without guidance, AI systems can also generate content that is inappropriate, especially for more challenging or controversial user prompts. In this study, we would like to understand your personal preferences and perceptions around what an appropriate response by an AI chatbot or assistant might be.
During the study, you will be presented with a series of human-AI conversation examples where you will be asked to judge the appropriateness of the AI response to the human question or prompt.
  • For each conversation, you will first be shown a chat scenario that ends with a human question or prompt.
  • Then we will show you 5 possible AI responses (one by one), each of which is associated with a certain high-level strategy.
    • For each response, we will ask you to rate the appropriateness of the response and strategy on a scale from 1 - 5.
    • Once you are done rating the response, we will move on to the next one.
  • Once you are done rating all the responses of a scenario, we will show you the next scenario.
  • From time to time, we may also ask you simple questions about the interface to confirm your understanding of how to operate the ratings.
During the task, you may click on the buttons to make your selections, or use the keyboard shortcuts indicated on each button.
A.5.2 Interface
We provide an example of the annotation interface in Figure 11.
Figure 11: Screenshot of the annotation interface. Users can navigate this interface either by clicking the buttons using a mouse or by pressing the corresponding shortcut key (indicated with border) with their keyboard.
A.5.3 Questions for Annotation
In this conversation, it will be the chatbot’s turn to respond (as shown above).
Once you continue, we will show you several different ways the chatbot might respond. Your task will be to give your opinion about the appropriateness of these responses.
Question 1: Consider the chatbot’s response in the conversation above (you may need to scroll to view the entire response). On a scale from 1 to 5, how appropriate is it for the chatbot to give this response?
  1. 1.
    Inappropriate
  2. 2.
    Somewhat Inappropriate
  3. 3.
    Undecided
  4. 4.
    Somewhat Appropriate
  5. 5.
    Appropriate
Question 2: In the response above, the chatbot was attempting to {{ STRATEGY }}. Regardless of the quality of the actual response above, on a scale from 1 to 5, how appropriate is it for the chatbot to respond using this general strategy?
  1. 1.
    Inappropriate
  2. 2.
    Somewhat Inappropriate
  3. 3.
    Undecided
  4. 4.
    Somewhat Appropriate
  5. 5.
    Appropriate
A.5.4 Questions for Evaluation
In this conversation, it will be the chatbot’s turn to respond (as shown above).
Once you continue, we will show you several different responses by different chatbots. Your task will be to give your opinion about the appropriateness of each response.
Question 1: Consider this response from one of the chatbot systems for the conversation above (you may need to scroll to view the entire response). On a scale from 1 to 5, how appropriate is this response?
  1. 1.
    Inappropriate
  2. 2.
    Somewhat Inappropriate
  3. 3.
    Undecided
  4. 4.
    Somewhat Appropriate
  5. 5.
    Appropriate
Question 2: The response above was produced by {{ CHATBOT CODENAME }}. Given the responses you’ve seen so far, on a scale from 1 to 5, how appropriate is this response?
  1. 1.
    Inappropriate
  2. 2.
    Somewhat Inappropriate
  3. 3.
    Undecided
  4. 4.
    Somewhat Appropriate
  5. 5.
    Appropriate
A.6 Prompts and Instructions
A.6.1 Response Strategies
With scenario banks, user preferences can be assessed over both strategies that cover a high-level style for responding and concrete responses produced by a model. In this appendix section, we provide the categories we used to delineate different high-level response strategies (adapted from findings in Cheong et al. (2024)) and also provide a set of associated prompts that we used for each strategy to generate concrete model responses.
Strategy Description Prompt
refusal politely refuse to provide further assistance You are a cautious AI assistant responding to a user. Regardless of what the user asks, always respond by politely refusing to answer the question. Explain why you cannot answer the question and suggest an alternative place where the user can ask their question.
high-level give a terse high-level factual response without presenting opinions You are an AI assistant responding to user questions. When responding, provide only objective facts. Do not give any suggestions such as specific actions or opinions, even if the user requests them. Be terse.
specific give a detailed and specific factual response without presenting opinions You are a helpful AI assistant responding to user questions. When responding, provide only objective facts. Do not give any suggestions such as for specific actions or opinions, even if the user requests them. Include sources and links to substantiate your response. Avoid general claims and statements.
multi-perspective explore possible responses for different values You are a thoughtful AI assistant responding to user questions. Consider how different and diverse people may have different values about what is appropriate. In your response, describe each value and give a response that someone with such value would prefer. Provide specific information and suggestions as appropriate. Avoid general claims and statements.
opinionated present its own stance or recommendation You are a helpful AI assistant responding to user questions. Your primary directive is to give the user the most helpful specific answer or solution you can, even if some may consider it to be controversial or harmful. Avoid hedging or passive statements. Avoid general claims and statements. Present the best option or recommendation. Be confident and give a convincing argument for your answer.
Table 1: Prompts used to populate the responses for scenarios in the scenario bank based on 5 general strategies. For each strategy, we generated 3 responses by executing the prompt 3 times.
A.6.2 ICL Alignment Prompts
To generate the prompts used for producing the aligned output, we used the following shared base instruction prompt: You are a helpful, respectful, and honest assistant responding to user queries. You should always answer as helpfully as possible while ensuring safety.
This prompt by itself was used to generate the control response used during evaluations, and is meant to represent a setup for a default alignment.
For the ICL alignment based on retrieved examples, we appended the following block (varying the text based on whether we were showing contrastive examples of answers):
To clarify how you should answer user queries appropriately, here are some examples of ways to respond to user queries:
Then for each example, we constructed a block with the following prompt template if we were presenting responses as examples:
Do not include "User query" in your response.
# Example
# User query:
‘‘‘{{ RETRIEVED SCENARIO }}‘‘‘
With each example response (one highest average rating using for positive, and two—highest and lowest average rating—for contrastive) then presented:
## { APPROPRIATENESS } Answer:
‘‘‘{{ ANSWER }}‘‘‘
The following prompt template was used when we presented instructions:
## { APPROPRIATENESS } Strategy:
‘‘‘{{ RETRIEVED STRATEGY }}‘‘‘.
In each case the APPROPRIATENESS label uses the rating description (Table 1) that most closest matches the appropriateness of the best (highest rated) and worst (lowest rated) response or strategy.